# Multi-Treatment-DML: Causal Estimation for Multi-Dimensional Continuous Treatments with Monotonicity Constraints in Personal Loan Risk Optimization

### v1

@@ 3 Preliminary Propositions
[…]
@@ 4.1 Propensity Network
The Propensity Network predicts the treatment propensity $\hat{\mathbf{T}}=\mathbf{E}[\mathbf{T}|\mathbf{X}]$, the outcome $\hat{\textit{Y}}=\mathbf{E}[\textit{Y}|\textbf{X}]$, and a sensitivity coefficient $K$—which is used in the causal network to enforce monotonicity between $\Delta\textbf{T}$ and $\Delta\textit{Y}$— from the observed covariates $X$.
@@
In our experiments, we set the Tweedie power parameter $\rho$ to 1.91.91.9, which is commonly used for modeling zero-inflated, long-tailed financial outcomes. Weight Orthogonality Regularization: We introduce a criterion called Representation Layer Orthogonality (RLO), which provides an intuitive approach to obtaining non-overlapping factors:
@@ 4.3 Overall Loss Function
$\mathcal{L}_{\text{total}}=$ $\mathcal{L}_{\text{treatment}}+\mathcal{L}_{\text{outcome}}+\mathcal{L}_{\text{final}}$
@@ 5.3 Comparative Results on Publication Datasets
In Table 1, Multi-Treatment-DML achieves the best individual treatment effect estimation accuracy ($\sqrt{\epsilon_{\text{PEHE}}}$) and average treatment effect error ($\epsilon_{\text{ATE}}$) on the Twins dataset, with values of 0.323±0.0020.323$\pm$ 0.0020.323 ± 0.002 and 0.011±0.0010.011$\pm$ 0.0010.011 ± 0.001, respectively, significantly outperforming all baseline methods.
Table 2 presents a comparative evaluation of Multi-Treatment DML and baseline methods on the Jobs and Lazada datasets. For the Jobs dataset, Multi-Treatment DML achieves the lowest policy risk ($\mathcal{R}_{\mathrm{pol}}(\pi)=0.221$) and ATT error ($\epsilon_{\text{ATT}}=0.064$), outperforming all baselines. On the Lazada dataset, Multi-Treatment DML matches the best policy risk (0.9930.9930.993) and achieves competitive ATT error (0.0430.0430.043), while Qini AUUC (0.0180.0180.018) is comparable to the best baseline. Overall, Multi-Treatment-DML demonstrates superior or competitive results across all metrics, confirming its effectiveness in both policy optimization and treatment effect estimation. More importantly, our method is capable of handling arbitrary-dimensional and continuous treatments setting where existing methods are rarely applicable. This represents a significant advantage and the key innovation of our approach.
@@
T-Learner 0.341±0.0050.341$\pm$ 0.0050.341 ± 0.005 0.053±0.0120.053$\pm$ 0.0120.053 ± 0.012
S-Learner 0.340±0.0020.340$\pm$ 0.0020.340 ± 0.002 0.047±0.0090.047$\pm$ 0.0090.047 ± 0.009
X-Learner 0.338±0.0030.338$\pm$ 0.0030.338 ± 0.003 0.056±0.0120.056$\pm$ 0.0120.056 ± 0.012
TARNet 0.336±0.0020.336$\pm$ 0.0020.336 ± 0.002 0.033±0.0090.033$\pm$ 0.0090.033 ± 0.009
DragonNet 0.338±0.0030.338$\pm$ 0.0030.338 ± 0.003 0.035±0.0090.035$\pm$ 0.0090.035 ± 0.009
EFIN 0.338±0.0050.338$\pm$ 0.0050.338 ± 0.005 0.026±0.0040.026$\pm$ 0.0040.026 ± 0.004
@@
T-Learner 0.389±0.0600.389$\pm$ 0.0600.389 ± 0.060 0.105±0.0340.105$\pm$ 0.0340.105 ± 0.034 0.992±0.0000.992$\pm$ 0.0000.992 ± 0.000 0.042±0.0000.042$\pm$ 0.0000.042 ± 0.000 0.002±0.0000.002$\pm$ 0.0000.002 ± 0.000
S-Learner 0.345±0.0680.345$\pm$ 0.0680.345 ± 0.068 0.117±0.0410.117$\pm$ 0.0410.117 ± 0.041 $\mathbf{0.990\pm 0.000}$ 0.042±0.0000.042$\pm$ 0.0000.042 ± 0.000 $\mathbf{0.099\pm 0.000}$
X-Learner 0.565±0.1100.565$\pm$ 0.1100.565 ± 0.110 0.084±0.0370.084$\pm$ 0.0370.084 ± 0.037 0.985±0.0000.985$\pm$ 0.0000.985 ± 0.000 0.043±0.0000.043$\pm$ 0.0000.043 ± 0.000 0.054±0.0010.054$\pm$ 0.0010.054 ± 0.001
TARNet 0.479±0.1080.479$\pm$ 0.1080.479 ± 0.108 0.180±0.0520.180$\pm$ 0.0520.180 ± 0.052 0.987±0.0010.987$\pm$ 0.0010.987 ± 0.001 0.075±0.0060.075$\pm$ 0.0060.075 ± 0.006 0.002±0.0000.002$\pm$ 0.0000.002 ± 0.000
DragonNet 0.768±0.0980.768$\pm$ 0.0980.768 ± 0.098 0.192±0.0310.192$\pm$ 0.0310.192 ± 0.031 0.987±0.0020.987$\pm$ 0.0020.987 ± 0.002 0.071±0.0090.071$\pm$ 0.0090.071 ± 0.009 0.018±0.0270.018$\pm$ 0.0270.018 ± 0.027
EFIN 0.294±0.0450.294$\pm$ 0.0450.294 ± 0.045 0.107±0.0380.107$\pm$ 0.0380.107 ± 0.038 0.993±0.0000.993$\pm$ 0.0000.993 ± 0.000 $\mathbf{0.017\pm 0.002}$ 0.002±0.0000.002$\pm$ 0.0000.002 ± 0.000
MTDML $\mathbf{0.221\pm 0.009}$ $\mathbf{0.064\pm 0.024}$ $\mathbf{0.993\pm 0.000}$ 0.043±0.0010.043$\pm$ 0.0010.043 ± 0.001 0.018±0.0010.018$\pm$ 0.0010.018 ± 0.001
@@ 5.4 Overall Performance on Personal Loan Dataset
For evaluation, we adopt the PCOC (Predict Click Over Click) metric. PCOC (Predict Click Over Click) is a commonly used metric in online advertising to evaluate the calibration of predicted click-through rates (CTR). It is defined as the ratio between the average predicted CTR and the average observed (empirical) CTR: A PCOC value close to 1 indicates well-calibrated predictions, and values less than 1 indicate underestimation. As shown in Table 3, the best results, with PCOC closest to 111 (1.0021.0021.002) and the highest Qini AUUC (0.1870.1870.187). These results indicate that the proposed enhancements lead to more accurate and better-calibrated uplift modeling, with Qini AUUC increasing and PCOC approaching 111, which is optimal.
@@
Table 4 presents the comparative results of various methods on the Personal Loan dataset in terms of PCOC and Qini AUUC metrics. As shown, Multi-Treatment-DML achieves the highest Qini AUUC score of 0.2430.2430.243, significantly outperforming all baseline methods, indicating its superior uplift modeling and ranking performance. Meanwhile, Multi-Treatment-DML’s PCOC is 1.1001.1001.100, which is closer to the ideal calibration value of 111, demonstrating better consistency between predicted and actual click-through rates. Other methods, such as DragonNet and TarNet, also perform well on Qini AUUC, but their PCOC values deviate more from 111, suggesting some degree of underestimation. Overall, Multi-Treatment-DML balances prediction accuracy and ranking capability in the personal loan scenario, validating the effectiveness and business value of the proposed method.
@@
S-Learner 0.8250.8250.825 0.2140.2140.214
T-Learner 0.8450.8450.845 0.1220.1220.122
X-Learner 0.8300.8300.830 0.1560.1560.156
TARNet 0.8910.8910.891 0.2030.2030.203
DragonNet 0.8850.8850.885 0.1970.1970.197
EFIN 0.8260.8260.826 0.0600.0600.060

### v2

@@ 3 Preliminary Propositions
As illustrated in Figure 1, $\textbf{X}\in\mathbb{R}^{D}$ denotes the observed covariates—user characteristics and behavioral features available prior to intervention. In the context of personal loan, X includes fundamental attributes and historical behaviors, such as actual cost, overdue counts, total repayments, and other relevant user information. The treatment vector $\textbf{T}\in\mathbb{R}^{K}$ represents $K$-dimensional continuous financial offer variables (e.g., credit limit, interest rate, loan term with $K=3$), which directly affect the outcome $\textit{Y}\in\mathbb{R}$. Unlike binary or categorical interventions ($K=1$, $\textit{T}\in\{0,1\}$), financial treatments are inherently multi-dimensional and continuous. The outcome variable Y denotes the observed response, such as utilized loan amount. It is important to emphasize that our dataset consists solely of observational data, where treatment assignments are determined by internal business rules rather than randomized protocols.
@@ 4.1 Propensity Network
The Propensity Network predicts the treatment propensity $\hat{\mathbf{T}}=\mathbf{E}[\mathbf{T}|\mathbf{X}]$, the outcome $\hat{\textit{Y}}=\mathbf{E}[\textit{Y}|\textbf{X}]$, and a sensitivity coefficient KK—which is used in the causal network to enforce monotonicity between $\Delta\textbf{T}$ and $\Delta\textit{Y}$— from the observed covariates $X$.
@@
In our experiments, we set the Tweedie power parameter $\rho$ to $1.9$, which is commonly used for modeling zero-inflated, long-tailed financial outcomes. Weight Orthogonality Regularization: We introduce a criterion called Representation Layer Orthogonality (RLO), which provides an intuitive approach to obtaining non-overlapping factors:
@@ 4.3 Overall Loss Function
$\mathcal{L}_{\text{total}}=\$$\mathcal{L}_{\text{treatment}}+\mathcal{L}_{\text{outcome}}+\mathcal{L}_{\text{final}}$
@@ 5.3 Comparative Results on Publication Datasets
In Table 1, Multi-Treatment-DML achieves the best individual treatment effect estimation accuracy ($\sqrt{\epsilon_{\text{PEHE}}}$) and average treatment effect error ($\epsilon_{\text{ATE}}$) on the Twins dataset, with values of $0.323\pm 0.002$ and $0.011\pm 0.001$, respectively, significantly outperforming all baseline methods.
Table 2 presents a comparative evaluation of Multi-Treatment DML and baseline methods on the Jobs and Lazada datasets. For the Jobs dataset, Multi-Treatment DML achieves the lowest policy risk ($\mathcal{R}_{\mathrm{pol}}(\pi)=0.221$) and ATT error ($\epsilon_{\text{ATT}}=0.064$), outperforming all baselines. On the Lazada dataset, Multi-Treatment DML matches the best policy risk ($0.993$) and achieves competitive ATT error ($0.043$), while Qini AUUC ($0.018$) is comparable to the best baseline. Overall, Multi-Treatment-DML demonstrates superior or competitive results across all metrics, confirming its effectiveness in both policy optimization and treatment effect estimation. More importantly, our method is capable of handling arbitrary-dimensional and continuous treatments setting where existing methods are rarely applicable. This represents a significant advantage and the key innovation of our approach.
@@
T-Learner $341\pm 0.005 0.0$$053\pm 0.012$
S-Learner $340\pm 0.002 0.0$$047\pm 0.009$
X-Learner $338\pm 0.003 0.0$$056\pm 0.012$
[…]
DragonNet $338\pm 0.003 0.0$$035\pm 0.009$
EFIN $338\pm 0.005 0.0$$026\pm 0.004$
@@
T-Learner $0.389\pm 0.060$ $0.105\pm 0.034$ $992\pm 0.000 0.0$$42\pm 0.000 0.0$$002\pm 0.000$
[…]
X-Learner $565\pm 0.110 0.0$$084\pm 0.037$ $985\pm 0.000 0.0$$43\pm 0.000 0.0$$054\pm 0.001$
TARNet $0.479\pm 0.108$ $0.180\pm 0.052$ $987\pm 0.001 0.0$$75\pm 0.006 0.0$$002\pm 0.000$
DragonNet $0.768\pm 0.098$ $0.192\pm 0.031$ $987\pm 0.002 0.0$$71\pm 0.009 0.0$$018\pm 0.027$
[…]
MTDML $\mathbf{0.221\pm 0.009}$ $\mathbf{0.064\pm 0.024}$ $\mathbf{0.993\pm 0.000}$ $043\pm 0.001 0.0$$018\pm 0.001$
@@ 5.4 Overall Performance on Personal Loan Dataset
For evaluation, we adopt the PCOC (Predict Click Over Click) metric. PCOC (Predict Click Over Click) is a commonly used metric in online advertising to evaluate the calibration of predicted click-through rates (CTR). It is defined as the ratio between the average predicted CTR and the average observed (empirical) CTR: A PCOC value close to 1 indicates well-calibrated predictions, and values less than 1 indicate underestimation. As shown in Table 3, the best results, with PCOC closest to 11 ($1.002$) and the highest Qini AUUC ($0.187$). These results indicate that the proposed enhancements lead to more accurate and better-calibrated uplift modeling, with Qini AUUC increasing and PCOC approaching $1$, which is optimal.
@@
Table 4 presents the comparative results of various methods on the Personal Loan dataset in terms of PCOC and Qini AUUC metrics. As shown, Multi-Treatment-DML achieves the highest Qini AUUC score of $0.243$, significantly outperforming all baseline methods, indicating its superior uplift modeling and ranking performance. Meanwhile, Multi-Treatment-DML’s PCOC is $1.100$, which is closer to the ideal calibration value of $1$, demonstrating better consistency between predicted and actual click-through rates. Other methods, such as DragonNet and TarNet, also perform well on Qini AUUC, but their PCOC values deviate more from $1$, suggesting some degree of underestimation. Overall, Multi-Treatment-DML balances prediction accuracy and ranking capability in the personal loan scenario, validating the effectiveness and business value of the proposed method.
@@
S-Learner $0.825$ $0.214$
T-Learner $0.845$ $0.122$
X-Learner $0.830$ $0.156$
TARNet $0.891$ $0.203$
DragonNet $0.885$ $0.197$
EFIN $0.826$ $0.060$